OpenStudy (anonymous):

Can someone show me the steps for this? (calc). The velocity of a particle moving on a line is v=ds/dt= 6 sin (2t) m/sec for all t. If s=0 when t=0, find s when t= pi/2 secs.

OpenStudy (anonymous):

@johnny101 v = ds/dt can we write ds = vdt ?????

OpenStudy (anonymous):

ds = 6 sin(2t) dt is it possible for u do the integration of ds?

OpenStudy (anonymous):

by using pi/2 as the upper and 0 as the lower?

OpenStudy (anonymous):

no need to apply limits, first do the integration as an indefinite integral we will use the boundary condition after determining s

OpenStudy (anonymous):

can you help me with that im sorry im rusty on integrals

OpenStudy (anonymous):

\[s =6* \frac{ \cos(2t) }{ 2 } + constant\] first boundary condition is s = 0 when t = 0 , hope u can determine the constant now

OpenStudy (anonymous):

so for that setup you just did the deriv of sin? or did you use x^n+1/n +1? trying to follow the reasoning here..what you have is just 0

OpenStudy (anonymous):

i did the integration of sinx it is -cosx i forgot to put the negative sign s = 0 when t=0 \[0 = -6 * \frac{ \cos(0) }{ 2 } + constant\] 0 = -3 +constant as cos(0) = 1 therefore ,constant = 3 hence \[s= -6*\frac{ \cos(2t) }{ 2 } + 3\] now just put t = pi/2

OpenStudy (anonymous):

did u follow the steps?

OpenStudy (anonymous):

ok so sin -> -cos x i got that, where did the formation come from of -6* cos(0)/2 +C?

OpenStudy (anonymous):

we get \[s = -6*\frac{ \cos(2t) }{ 2 } + constant\] when we solve an indefinite integral we always get some constant term now the condition is given s = 0 when t = 0 i just plug in the value of s and t which is 0

OpenStudy (anonymous):

right, im saying, we have 6 sin 2t right? so you integrated which made sin -> -cos, which you then applied to the 6 right? so -6 * cos(2t), where is the / 2 coming from? from basic integration x^n+1/n+1? wouldn't that give 3? lost on the denominator i know its somethhing basic.

OpenStudy (anonymous):

2 in the denominator has come from the chain rule of integration

OpenStudy (anonymous):

can you show that please?

OpenStudy (anonymous):

\[s = -6*\frac{ \cos(2t) }{ \frac{ d }{ dt }(2t) } + constant\] \[s = -6*\frac{ \cos(2t) }{ 2 } + constant\]

OpenStudy (anonymous):

i have the chain rule in my notes involving taking the exponent to the front, subtracting the exponent by one and multiplying by the deriv of what's inside... so on this we have 6sin 2t. How did you convert that to 2t in fraction form? can you list it out(im sorry not grasping it from that directly)

OpenStudy (anonymous):

this is the way of solving an integral using chain rule we always take the derivative and it should be in the denominator this is the rule , mate

OpenStudy (anonymous):

so 2t is just the derivative of 6 sin 2t which is where that came from?

OpenStudy (anonymous):

derivative in the denominator is a rule u have to remember it for example u need to determine integral of sin(5t+1) we will get a solution like this \[\frac{ -\cos(5t+1) }{ \frac{ d }{ dt }(5t+1)}\] \[\frac{ -\cos(5t+1) }{ 5 }\]

OpenStudy (anonymous):

ok so, 6 sin2t we integrate that which makes it -cos for the sin, with the negative transferring to the 6, so -6* cos 2t, whose deriv is 2t, so -6*cos2t/2t

OpenStudy (anonymous):

\[-6*\frac{ \cos(2t) }{ \frac{ d }{ dt }(2t) }\] u need to take the derivative of 2t

OpenStudy (anonymous):

so we just take the 2t and put it in both the numerator and denominator, the deriv of 2t is just 2, which is why i wasnt following

myininaya (myininaya):

\[ \text{ Let } a \text{ and } b \text{ be constants }\] \[\int\limits_{}^{}b \sin(at ) dt \text{ where } a \neq 0\] Let us do a substitution \[at=u => a dt =du \] So we have in terms of u the integral is: \[\int\limits_{}^{} b \sin(u) \frac{1}{a} du \text{ since } dt=\frac{1}{a} du\] \[\text{ now } \frac{b}{a} \text{ is a constant } \] so we have \[\frac{b}{a} \int\limits_{}^{}\sin(u) du=\frac{b}{a} (-\cos(u))+C =\frac{-b}{a}\cos(u)+C\] So the final answer in terms of t is \[\frac{-b}{a} \cos(at)+C \text{ since } u=at \] Now this is what I assumed you were pm-ing me about...Maybe.

OpenStudy (anonymous):

well i was referring to the setup of it but yes. so based off what we were using before, that gives us with -6*cos 2t/2 + C,

OpenStudy (anonymous):

after that i then..?

myininaya (myininaya):

it says when t=0 s=0 so they are giving you info to find your C

myininaya (myininaya):

once you find your C you have one more step

OpenStudy (anonymous):

if i just input t=0 into that that gives me the 0 back, so i input pi/2 into that? so -6*cos(2pi/2)/2+C?

myininaya (myininaya):

You said s(t)=-3 cos(2t)+C

myininaya (myininaya):

Now it tells you when s=0 t=0 use that to find C

myininaya (myininaya):

Then we can do the last step

myininaya (myininaya):

Replace t with 0 and replace s with 0 like this 0=-3cos(2*0)+C solve this equation for C

OpenStudy (anonymous):

C=cos no since the 0 multiplies out to just 0 and 0/-3 = 0

myininaya (myininaya):

so cos(0)=1 correct?

myininaya (myininaya):

maybe you can solve this equation 0=-3+C

OpenStudy (anonymous):

c =3, but how did you conclude cos(0)=1 if that solves to 0?

myininaya (myininaya):

unit circle

OpenStudy (anonymous):

ok roger

myininaya (myininaya):

So we have s=-3 cos(2t)+3

myininaya (myininaya):

now figure out what S is when t=pi/2

myininaya (myininaya):

do you know what cos(pi) is?

OpenStudy (anonymous):

-1